OpenStudy (anonymous):

select a solution to the equation y = 3x +1.

Directrix (directrix):

I assume we have no options this time. Pick any x, say 0, and compute the y that goes with it according to the dictates of the equations. y = 3(0) + 1= 0+1 = 1 Selected solution: (0,1).

OpenStudy (anonymous):

Oh shoot i'm sorry hold on

OpenStudy (anonymous):

A. (1,-4) B. (1,-5) C. (1,5) D. (1,4)

Directrix (directrix):

y = 3x +1 On first glance, it appears that D. (1,4) is the only solution. Let's see if it works: 4 = 3(1) + 1 4 = 4 So, D. (1,4) is a solution. Do you see any others?

OpenStudy (anonymous):

I believe D is the only one.

Directrix (directrix):

Me, too.